A researcher hypothesizes that people who listen to music via headphones have greater hearing loss and will thus score lower on a hearing test than those in the general population. On a standard hearing test, the overall mean for the general population is 22.5. The researcher gives this same test to a random sample of 12 individuals who regularly use headphones. Their scores on the test are 15, 14, 20, 20, 25, 22, 21, 19, 16, 17, 21, 22.
1
Expert's answer
2013-04-15T11:52:39-0400
a) Breakeven output will be in point, where TR = TC,
TR = 18Q, TC = 3,500,000 + 10Q
TR = TC
18Q = 3,500,000 + 10Q
8Q = 3,500,000
Q = 437,500
In dollars: 437,500*18 = $7,875,000 is the breakeven output.
b) for $1,500,000 profit:
TR = TC + 1,500,000
18Q = 5,000,000 + 10Q
8Q = 5,000,000
Q = 625,000 units
c) If Q = 400,000:
Leverage = 18*400,000 - 3,500,000 - 10*400,000 = -$300,000
d) The probability will be = (362,500 + 100,000 - 437,500)/100,000 = 25% for
incurring of loss

(d) Assuming that sales of oil are normally distributed with a mean of
362,500 barrels and a standard deviation of 100,000 barrels, determine the
probability that Offshore will incur an operating loss.